a) On a throw-in, a tennis player hits his 0.075 Kg ball with the racket, providing him with an impulsive force. Assuming that the ball and the racket are in contact for 5 ms and then the ball leaves the racket at a speed of 40 m / s, what would be approximately the average racket force on the ball? Ignore all other forces, including gravity

b) Water flows out of a hose at a rate of 1.2 Kg for every second with a speed of 15 m / s and goes to a wall that stops it. Any splash back is ignored. What is the average force exerted by the wall on water? Solution
Answer #1

a)

F = m ( v - 0) /.t

F = 0.075* 40 / 0.005 = 600 N

====

b)

F = m v / t = (m/t) v

F = 1.2* 15 = 18 N

=====

Comment before rate in case any doubt, will reply for sure... goodluck